List the first five terms of the sequence.
f(n+1)=2f(n)+n. Where f(1)=1, for n > or = 1

Answer by jim_thompson5910(35256)

First term: 1

This is given when they say f(1) = 1

-------------------------------------------------------

Second term

f(n+1)=2f(n)+n

f(1+1)=2f(1)+1

f(2)=2*1+1

f(2)=3

The second term is 3


-------------------------------------------------------

Third term

f(n+1)=2f(n)+n

f(2+1)=2f(2)+2

f(3)=2*3+2

f(3)=8

The third term is 8

-------------------------------------------------------

Fourth term

f(n+1)=2f(n)+n

f(3+1)=2f(3)+3

f(4)=2*8+3

f(4)=19

The fourth term is 19

-------------------------------------------------------

Fifth term

f(n+1)=2f(n)+n

f(4+1)=2f(4)+4

f(5)=2*19+4

f(5)=42

The fifth term is 42

-------------------------------------------------------

The first five terms are: 1, 3, 8, 19, 42